# Properties of Immersions for Systems with Multiple Limit Sets with Implications to Learning Koopman Embeddings

### v1

@@ 2.2 Technical Definitions
The forward implication is well known. We recall the standard proof here. Suppose $\varphi(t,\xi)$ is precompact in ${\mathcal{X}}$. Let $t_{n}$ be a sequence such that $t_{n}\rightarrow\infty$. By Definition 2.9, there exists a subsequence $t_{n_{k}}$ such that $\varphi(t_{n_{k}},\xi)$ converges to a point $x$ in the closure of $\{\varphi(t,\xi)\mid t\in\mathbb{R}_{\geq 0}\}$. Thus, $\omega^{+}(\xi)$ contains $x$ and is nonempty. Now suppose that the system in (1) is linear, that is, $\dot{x}=Ax$ for some $A\in\mathbb{R}^{n\times n}$. If a solution $\varphi(t,\xi)$ of the linear system is unbounded, it can be shown that $\lim_{t\to\infty}\|\varphi(t,\xi)\|_{2}=\infty$. Then, $\omega^{+}(\xi)$ is empty since for any sequence $t_{n}\rightarrow\infty$, $\|\varphi(t_{n},\xi)\|_{2}\rightarrow\infty$. Thus, if $\omega^{+}(\xi)$ is nonempty, $\varphi(t,\xi)$ is bounded and thus is precompact in $\mathbb{R}^{n}$. Since ${\mathcal{X}}$ is closed, the closure of $\varphi(t,\xi)$ in $\mathbb{R}^{n}$ is contained in ${\mathcal{X}}$, which implies that $\varphi(t,\xi)$ is precompact in ${\mathcal{X}}$. $\Box$
Definition 2.11
Let $\mathcal{W}^{+}$ be the set of all $\omega$-limit sets of $\dot{x}=f(x)$ in ${\mathcal{X}}$. For each $\Omega\in\mathcal{W}^{+}$, we define its domain of attraction by
@@
By definition, the set $\mathcal{W}^{+}$ contains all the equilibria or closed orbits in ${\mathcal{X}}$.
@@
Recall that $\mathcal{W}^{+}$ is the set of all $\omega$-limit sets in ${\mathcal{X}}$.
Definition 2.12
A system of the form (1) has closed basins if the domain of attraction $D^{+}(\Omega)$ is closed for all $\omega$-limit sets $\Omega\in\mathcal{W}^{+}$.
The following lemma provide sufficient conditions for systems to have closed basins, which are satisfied by all linear systems.
Lemma 2.13
@@
  • (C2)
[…]
@@
Let $t_{k}\rightarrow+\infty$ be a sequence such that $\varphi(t_{k},\xi_{1})\rightarrow p$. By (C1), the sequence $\varphi(t_{k},\xi_{2})$ is contained in a compact subset of $X$. Therefore, there exists a subsequence $t_{k}^{\prime}$ of $t_{k}$ such that $\varphi(t_{k}^{\prime},\xi_{2})\rightarrow q$ for some $q$ in $\omega^{+}(\xi_{2})$. By (C2),
@@
Now we pick any point $p\in\omega^{+}(x)$. By the claim, there exists a sequence $q_{k}$ in $\omega^{+}(x_{k})$ such that $\|q_{k}-p\|\rightarrow 0$. Since $\Omega$ is closed, it follows that $p\in\Omega$. Since $p$ is arbitrary, $\omega^{+}(x)$ is a subset of $\Omega$.
@@
Remark 2.14
If ${\mathcal{X}}$ is a closed subset of a finite-dimensional normed space, the condition (C1) in Lemma 2.13 can be replaced with the condition that for every $\Omega\in\mathcal{W}^{+}$, there exists one trajectory in $\mathcal{D}^{+}(\Omega)$ that is precompact in ${\mathcal{X}}$.
Corollary 2.15
@@
We want to prove the corollary by showing that any linear system satisfies the conditions (C1) and (C2) in Lemma 2.13. The condition (C1) holds for any linear systems according to Lemma 2.10.
@@ 3 Main Theorem
Combining Lemma 2.15 and Corollary 3.17, the following corollary states a necessary condition for the existence of one-to-one linear immersions (or in general one-to-one immersions to systems with closed basins).
Corollary 3.18
@@
For non-existence of linear immersions as in Corollary 3.18 both precompactness of trajectories and the existence of countable but more than one $\omega$-limit sets are not only sufficient conditions, they are indeed necessary in the following sense. Arathoon and Kvalheim (2023) provide an example of a two dimensional system with two isolated equilibria with some trajectories that are neither precompact nor backward precompact (cf. Section 5.2) that admits a linear immersion. Similarly, there are systems with uncountably many $\omega$-limit sets that admit linear immersions, simplest examples being diffeomorphisms of linear systems with a nontrivial subspace as their equilibria.
Next, we demonstrate the application of our results through examples. We first show several examples where a one-to-one (linear) immersion is constructed manually when ${\mathcal{X}}$ does not satisfy the conditions in Corollary 3.18, but these immersions become discontinuous or ill-defined when we slightly modify ${\mathcal{X}}$ to violate one of these conditions. Sequentially, we offer examples from the literature where the existence of a linear immersion is uncertain, but our results establish that continuous one-to-one linear immersions do not exist.
Example 3.19
@@
However, if we extend ${\mathcal{X}}$ by a point to ${\mathcal{X}}^{\prime}=(-\infty,1]$, the function $F$ in (22) is not an immersion anymore, since $F(1)$ is not defined. This observation is explained by Corollary 3.18: Since ${\mathcal{X}}^{\prime}$ contains two limit sets $\{-1\}$ and $\{+1\}$, and all trajectories in $(-\infty,1]$ are precompact, there does not exist a one-to-one linear immersion for the system on ${\mathcal{X}}^{\prime}$.
Example 3.20
@@
with $|y|\leq 1$. That is, the system in (23) on ${\mathcal{X}}$ is immersed in the system in (21) on ${\mathcal{Z}}=[-1,1]$. In this example, $\mathcal{W}^{+}$ has two elements, and all trajectories of $x$ in $\mathbb{R}$ are precompact, but a one-to-one immersion exists. By Theorem 3.16, this is possible only if the system $\dot{y}=y^{2}-1$ does not have closed basins. Indeed, the domain of attraction $D^{+}(\{-1\})$ of the system of $y$ on ${\mathcal{Z}}$ is $[-1,1)$, not a closed set.
Furthermore, by using Example 3.19, the system of $y$ on $[-1,1)$ can be immersed in $\dot{z}=-2z$ with the immersion in (22). Thus, $\dot{x}=\sin(x)$ on ${\mathcal{X}}^{\prime}=(0,\pi]$ is immersed in $\dot{z}=-2z$ with the one-to-one mapping
@@
If we extend $\mathcal{X}^{\prime}$ to ${\mathcal{X}}=[0,\pi]$, the function $F(x)$ in (25) is not defined at 00 and thus is not an immersion on the closed interval. This can be again explained by Corollary 3.18 since all the trajectories of $x$ are precompact, and the interval $[0,\pi]$ contains two limit sets.
Example 3.21
@@
The $\omega$-limit sets of the system are $\{-1\}$, $\{0\}$ and $\{1\}$. Let $-\0{\mathcal{X}}=\mathbb{R}\backslash\{0\}$. Define $y=x^{-2}-1$. Then, $\dot{y}$ satisfies
@@
Thus, the system in (26) on ${\mathcal{X}}$ is immersed in $\dot{y}=-2y$ with the immersion $F(x)=x^{-2}-1$. Similar to the previous example, ${\mathcal{X}}$ contains two $\omega$-limit sets, but each of its path-connected components contains only one $\omega$-limit set and thus the result is consistent with Corollary 3.18.
Example 3.22
Consider the two-dimensional system in Example 2.7. The $\omega$-limit sets of the system are the origin $\{0\}$ and the unit circle $\{x\mid\|x\|_{2}=1\}$. The linear immersion $F(x)$ in (5) is continuous in ${\mathcal{X}}$, but extending it to $\mathbb{R}^{2}$ makes $F(x)$ singular at the origin and thus no longer an immersion. This can be explained by Corollary 3.18: Since all trajectories of $x$ are precompact and the set $\mathbb{R}^{2}$ contains two $\omega$-limit sets, there does not exist a one-to-one linear immersion for the system of $x$ on $\mathbb{R}^{2}$.
Example 3.23
Example 2.8 shows that a discontinuous one-to-one linear immersion exists for the one-dimensional system in (15) with three isolated equilibria. We question if this system possesses a continuous one-to-one linear immersion over $\mathbb{R}$. This can be answered by Corollary 3.18: Since ${\mathcal{X}}$ contains three equilibria and all trajectories are precompact in ${\mathcal{X}}$, a continuous one-to-one linear immersion does not exist for this system.
Example 3.24
Consider the unforced Duffing system in Williams et al. (2015)
@@
This system has two asymptotically stable equilibria (±$(\pm 1,0)$ and one unstable equilibrium $(0,0)$, as shown by Fig. 3. Let ${\mathcal{X}}$ be the entire plane. Williams et al. (2015) suggests that this system does not admit a one-to-one linear immersion over ${\mathcal{X}}$, which is confirmed by our results. According to Corollary 3.18, since the system has three $\omega$-limit sets and all of its trajectories are precompact in ${\mathcal{X}}$, there does not exist a one-to-one linear immersion over ${\mathcal{X}}$.
@@
This system has two $\omega$-limit sets, namely an unstable equilibrium at the origin and a stable limit cycle, as shown in Fig. 4. Let ${\mathcal{X}}$ be the entire plane. Since all trajectories of the system are precompact in ${\mathcal{X}}$, by Corollary 3.18, there is no one-to-one linear immersion over ${\mathcal{X}}$.
@@
with $\sigma=10$, $b=8/3$, and $r=28$. According to Hirsch et al. (2012), there exists an invariant ellipsoid $\mathcal{E}$ centered at $(0,0,2r)$ that contains all the $\omega$-limit sets of the system, which include three equilibria and the Lorenz attractor. Let ${\mathcal{X}}=\mathcal{E}$. Since ${\mathcal{X}}$ is invariant and compact, all trajectories are precompact in ${\mathcal{X}}$. Thus, by Corollary 3.18, there does not exist a one-to-one linear immersion for the Lorenz system on ${\mathcal{X}}$.
@@ 4 Implication for Learning Linear Immersions from Data
Assumption 1
The system $\dot{x}=f(x)$ satisfies conditions (T2) and (T3) of Theorem 3.16 on a path-connected forward-invariant subset ${\mathcal{X}}$ of ${\mathcal{M}}$.
Generally, given a fixed sampling time $\tau$ and a set of $N$ pairs $\{(x_{l},x_{l}^{+})\}_{l=1}^{N}$ where $x_{l}^{+}=\varphi(\tau,x_{l})$ for all $l$, the task of learning linear immersions $F:{\mathcal{X}}\rightarrow{\mathcal{Z}}$ involves finding the following set
*$\mathcal{F}^{*}(\tau,N)=\{F^{*}\in\mathcal{C}({\mathcal{X}},{\mathcal{Z}})\mid\exists A^{*}\in\mathbb{R}^{m\times m},$
@@
where $\mathcal{C}({\mathcal{X}},{\mathcal{Z}})$ is the space of continuous functions from ${\mathcal{X}}$ to ${\mathcal{Z}}$. The state pairs $\{(x_{l},x_{l}^{+})\}_{l=1}^{N}$ can be extracted from a single trajectory or multiple trajectories of the system. Essentially, *$\mathcal{F}^{*}(\tau,N)$ is the set of continuous functions that satisfy the condition in (3) for linear immersions at finitely many points $\{x_{l}\}_{l=1}^{N}$ and a fixed time step $\tau$.
@@
$(F^{*},A^{*})\in\underset{\begin{subarray}{c}F\in\mathcal{C}({\mathcal{X}},{\mathcal{Z}}),\\ A\in\mathbb{R}^{m\times m}\end{subarray}}{\arg\min}$ $~{}\sum_{l=1}^{N}\|F(x_{l}^{\prime})-e^{A\tau}F(x_{l})\|$ (34)
@@
The set *$\mathcal{F}^{*}(\tau,N)$ in (31) corresponds to the solutions $F^{*}$ of this problem that give zero objective value, i.e., those interpolating the data perfectly.
Theorem 3.16 shows that under Assumption 1, any linear immersion $F$ satisfies that for all pairs $\Omega_{i}$, $\Omega_{j}$ in $\mathcal{W}^{+}$,
@@
A critical question here is if any learned linear immersion in *$\mathcal{F}^{*}(\tau,N)$ would also share this property in (36). As a sanity check, note that any constant function $F^{*}(\cdot)=C$ for some $C\in{\mathcal{Z}}$ belongs to *$\mathcal{F}^{*}(\tau,N)$ (with the corresponding $A^{*}=0$) and does map every $\omega$-limit set to the same set. However, this may not hold for every learned linear immersion in *$\mathcal{F}^{*}(\tau,N)$ since the learned linear immersion only satisfies (3) at finitely many points (finitely many constraints), while a true linear immersion needs to satisfy (3) everywhere in ${\mathcal{X}}$ at all times (uncountably many constraints). However, in what follows we show that any function that does not collapse all $\omega$-limit sets would be excluded from *$\mathcal{F}^{*}(\tau,N)$ for small enough $\tau$ and large enough $N$. The following theorem provides a crucial step for this result.
Theorem 4.27
Let $\{x_{l}\}_{l=1}^{\infty}$ be a dense subset of ${\mathcal{X}}$. Let $F$ be a continuous mapping from $\mathbb{R}^{n}$ to $\mathbb{R}^{m}$. If for all $t>0$, there exists a sampling time $\tau\in(0,t]$ such that $F\in\mathcal{F}^{*}(\tau,N)$ for all $N>1$, then $F$ is a linear immersion of the system.
@@
Suppose that a continuous function $F$ satisfies the conditions in Theorem 4.27. Then, for all $j\in\mathbb{N}$, there exists a sampling time $\tau_{j}\in(0,2^{-j}]$ such that $F\in\mathcal{F}^{*}(\tau_{j},N)$ for all $N\geq 1$. Clearly, the positive sequence $\tau_{j}$ converges to zero as $j$ goes to infinity. Also, for each $j$, $F\in\lim_{N\rightarrow\infty}\mathcal{F}^{*}(\tau_{j},N)=:\mathcal{F}^{*}(\tau_{j})$. Note that the limit *$\mathcal{F}^{*}(\tau_{j})$ is well defined since the set sequence *$\mathcal{F}^{*}(\tau_{j},N)$ monotonically shrinks as $N$ increases.
@@
We pick an arbitrary $t>0$. Since $\tau_{j}$ is positive and converging to zero, there exists $k_{j}\in\mathbb{N}$ such that the sequence $\bar{t}_{j}:=\sum_{i=1}^{j}k_{j}\cdot\tau_{j}$ converges to $t$ as $j$ goes to infinity. Note that for all $j$, (38) implies that
@@
Corollary 4.28
Let $\{x_{l}\}_{l=1}^{\infty}$ be a dense subset of ${\mathcal{X}}$. Let $F$ be a continuous function such that $F(\Omega_{1})\not=F(\Omega_{2})$ for some $\Omega_{1}$ and $\Omega_{2}\in\mathcal{W}^{+}$. Then, there exists $t^{*}>0$ such that for all sampling times $\tau\in(0,t^{*}]$, $F\not\in\mathcal{F}^{*}(\tau,N)$ for some $N>0$.
@@
Let $F$ be a continuous function that does not map every $\Omega\in\mathcal{W}^{+}$ to the same subset of $\mathbb{R}^{m}$ According to Theorem 3.16, since (T2) and (T3) hold, $F$ is not a linear immersion of the system. By the contrapositive of Theorem 4.27, for any $F$ not a linear immersion, there exists $t>0$ such that for all $\tau\in(0,t]$ and for some large enough $N$, $F$ is not in *$\mathcal{F}^{*}(\tau,N)$. $\Box$
Corollary 4.28 reveals that any immersion candidate $F$ that can distinguish at least two $\omega$-limit sets in ${\mathcal{X}}$ would always be ruled out from $\mathcal{F}(\tau,N)$ by a small enough sampling time $\tau$ and a large enough sample size $N$.
Remark 4.29
@@
$\limsup_{j\rightarrow\infty}\lim_{N\rightarrow\infty}\mathcal{F}^{*}(\tau_{j},N)\subseteq\mathcal{C}(\mathcal{W}^{+}).$ (41)
Remark 4.30

### v5

@@ 2.2 Technical Definitions
The forward implication is well known. We recall the standard proof here. Suppose $\varphi(t,\xi)$ is precompact in $\mathcal{X}$. Let $t_{n}$ be a sequence such that $t_{n}\rightarrow\infty$. By Definition 2, there exists a subsequence $t_{n_{k}}$ such that $\varphi(t_{n_{k}},\xi)$ converges to a point $x$ in the closure of $\{\varphi(t,\xi)\mid t\in\mathbb{R}_{\geq 0}\}$. Thus, $\omega^{+}(\xi)$ contains $x$ and is nonempty.
Now suppose that the system in (1) is linear, that is, $\dot{x}=Ax$ for some $A\in\mathbb{R}^{n\times n}$. If a solution $\varphi(t,\xi)$ of the linear system is unbounded, it can be shown that for all $M>0$, there exists $t_{M}\geq 0$ such that $\|\varphi(t,\xi)\|_{2}\geq M$ for all $t\geq t_{M}$. Then, $\omega^{+}(\xi)$ is empty since for any sequence $t_{n}\rightarrow\infty$, $\|\varphi(t_{n},\xi)\|_{2}\rightarrow\infty$. Thus, if $\omega^{+}(\xi)$ is nonempty, $\varphi(t,\xi)$ is bounded and thus is precompact in $\mathbb{R}^{n}$. Since $\mathcal{X}$ is closed, the closure of $\varphi(t,\xi)$ in $\mathbb{R}^{n}$ is contained in $\mathcal{X}$, which implies that $\varphi(t,\xi)$ is precompact in $\mathcal{X}$. $\Box$
Definition 3.
For any subset $\Omega$ of $\mathcal{X}$, we define its domain of attraction by
@@
By (13), if $\Omega$ is not an $\omega$-limit set, $D^{+}(\Omega)$ is empty.
@@
Definition 4.
Let $\mathcal{W}^{+}$ be the set of all nonempty $\omega$-limit sets of $\dot{x}=f(x)$ in $\mathcal{X}$. A system of the form (1) has closed basins if the domain of attraction $D^{+}(\Omega)$ is closed for all $\omega$-limit sets $\Omega\in\mathcal{W}^{+}$.
The following lemma provides sufficient conditions for systems to have closed basins, which are satisfied by all finite-dimensional linear systems.
Lemma 2.
@@
  2. (C2)
    The system is incrementally stable in the closure of $\mathcal{D}^{+}(\Omega)$. That is, there exists a function $\alpha$ of class $\mathcal{K}$ such that, for any two initial states $\xi_{1}$ and $\xi_{2}$ in $\emph{cl}(\mathcal{D}^{+}(\Omega))$ and for all $t\geq 0$, $\|\varphi(t,\xi_{1})-\varphi(t,\xi_{2})\|\leq\alpha(\|\xi_{1}-\xi_{2}\|)$.
@@
Let $t_{k}\rightarrow+\infty$ be a sequence such that $\varphi(t_{k},\xi_{1})\rightarrow p$. By (C1), the sequence $\varphi(t_{k},\xi_{2})$ is contained in a compact subset of $\mathcal{X}$. Therefore, there exists a subsequence $t_{k}^{\prime}$ of $t_{k}$ such that $\varphi(t_{k}^{\prime},\xi_{2})\rightarrow q$ for some $q$ in $\omega^{+}(\xi_{2})$. By (C2),
@@
Now we pick any point $p\in\omega^{+}(x)$. By the claim, there exists a sequence $q_{k}$ in $\omega^{+}(x_{k})$ such that $\|q_{k}-p\|\rightarrow 0$. Since $\Omega$ is closed [1], it follows that $p\in\Omega$. Since $p$ is arbitrary, $\omega^{+}(x)$ is a subset of $\Omega$.
@@
Remark 6.
If $\mathcal{X}$ is a closed subset of a finite-dimensional normed space, the condition (C1) in Lemma 2 can be replaced with the condition that for every $\Omega\in\mathcal{W}^{+}$, there exists one trajectory in $\mathcal{D}^{+}(\Omega)$ that is precompact in $\mathcal{X}$. We also note that we can define an abstract dynamical system over a metric space, and Lemma 2 can be generalized accordingly.
Corollary 1.
@@
We want to prove the corollary by showing that any linear system satisfies the conditions (C1) and (C2) in Lemma 2. The condition (C1) holds for any linear system according to Lemma 1.
@@ 3 Main Theorem
Combining Corollaries 1 and 2, the following corollary states a necessary condition for the existence of one-to-one linear immersions (or in general one-to-one immersions to systems with closed basins).
Corollary 3.
@@
For non-existence of linear immersions as in Corollary 3 both precompactness of trajectories and the existence of countable but more than one $\omega$-limit sets are not only sufficient conditions, they are indeed necessary in the following sense. The paper [2] provides an example of a two-dimensional system with two isolated equilibria with some trajectories that are neither precompact nor backward precompact (cf. Section 5.2) that admits a linear immersion. Similarly, there are systems with uncountably many $\omega$-limit sets that admit linear immersions, simplest examples being diffeomorphisms of linear systems with a nontrivial subspace as their equilibria.
Next, we demonstrate the application of our results through examples. We first show several examples where a one-to-one (linear) immersion is constructed manually when $\mathcal{X}$ does not satisfy the conditions in Corollary 3, but these immersions become discontinuous or ill-defined when we slightly modify $\mathcal{X}$ to violate one of these conditions. Sequentially, we offer examples from the literature where the existence of a linear immersion is unclear, but our results establish that continuous one-to-one linear immersions do not exist.
Example 3.
@@
However, if we extend $\mathcal{X}$ by a point to $\mathcal{X}^{\prime}=(-\infty,1]$, the function $F$ in (16) is not an immersion anymore, since $F(1)$ is not defined. This observation is explained by Corollary 3: Since $\mathcal{X}^{\prime}$ contains two $\omega$-limit sets $\{-1\}$ and $\{+1\}$, and all trajectories in $(-\infty,1]$ are precompact, there does not exist a one-to-one linear immersion for the system on $\mathcal{X}^{\prime}$.
Example 4.
@@
with $|y|\leq 1$. That is, the system in (17) on $\mathcal{X}$ is immersed in the system in (15) on $\mathcal{Z}=[-1,1]$. In this example, $\mathcal{W}^{+}$ has two elements, and all trajectories of $x$ in $\mathbb{R}$ are precompact, but a one-to-one immersion exists. By Theorem 1, this is possible only if the system $\dot{y}=y^{2}-1$ does not have closed basins. Indeed, the domain of attraction $D^{+}(\{-1\})$ of the system of $y$ on $\mathcal{Z}$ is $[-1,1)$, not a closed set.
Furthermore, by Example 3, the system of $y$ on $[-1,1)$ can be immersed in $\dot{z}=-2z$ with the immersion in (16). Thus, $\dot{x}=\sin(x)$ on $\mathcal{X}^{\prime}=(0,\pi]$ is immersed in $\dot{z}=-2z$ with the one-to-one mapping
@@
If we extend $\mathcal{X}^{\prime}$ to $\mathcal{X}=[0,\pi]$, the function $F(x)$ in (19) is not defined at 00 and thus is not an immersion on the closed interval. This can be again explained by Corollary 3 since all the trajectories of $x$ are precompact, and the interval $[0,\pi]$ contains two limit sets.
Example 5.
@@
The $\omega$-limit sets of the system are $\{-1\}$, $\{0\}$ and $\{1\}$. Let $\0\mathcal{X}=\mathbb{R}\backslash\{0\}$. Define $y=x^{-2}-1$. Then, $\dot{y}$ satisfies
@@
Thus, the system in (20) on $\mathcal{X}$ is immersed in $\dot{y}=-2y$ with the immersion $F(x)=x^{-2}-1$. Similar to the previous example, $\mathcal{X}$ contains two $\omega$-limit sets, but each of its path-connected components contains only one $\omega$-limit set and thus this observation is consistent with Corollary 3.
Example 6.
Consider the two-dimensional system in Example 1. The $\omega$-limit sets of the system are the origin $\{0\}$ and the unit circle $\{x\mid\|x\|_{2}=1\}$. The linear immersion $F(x)$ in (5) is continuous in $\mathcal{X}$, but extending it to $\mathbb{R}^{2}$ makes $F(x)$ singular at the origin and thus no longer an immersion. This can be explained by Corollary 3: Since all trajectories of $x$ are precompact and the set $\mathbb{R}^{2}$ contains two $\omega$-limit sets, there does not exist a one-to-one linear immersion for the system of $x$ on $\mathbb{R}^{2}$.
Example 7.
Example 2 shows that a discontinuous one-to-one linear immersion exists for the one-dimensional system in (9) with three isolated equilibria. We question if this system possesses a continuous one-to-one linear immersion over $\mathbb{R}$. This can be answered by Corollary 3: Since $\mathcal{X}$ contains three equilibria and all trajectories are precompact in $\mathcal{X}$, a continuous one-to-one linear immersion does not exist for this system.
Example 8.
Consider the unforced Duffing system in [35]
@@
This system has two asymptotically stable equilibria (±$(\pm 1,0)$ and one unstable equilibrium $(0,0)$, as shown by Fig. 4. Let $\mathcal{X}$ be the entire plane. The work in [35] suggests that this system does not admit a one-to-one linear immersion over $\mathcal{X}$, which is confirmed by our results. According to Corollary 3, since the system has three $\omega$-limit sets and all of its trajectories are precompact in $\mathcal{X}$, there does not exist a one-to-one linear immersion over $\mathcal{X}$.
@@
This system has two $\omega$-limit sets, namely an unstable equilibrium at the origin and a stable limit cycle, as shown in Fig. 5. Let $\mathcal{X}$ be the entire plane. Since all trajectories of the system are precompact in $\mathcal{X}$, by Corollary 3, there is no one-to-one linear immersion over $\mathcal{X}$.
@@
with $\sigma=10$, $b=8/3$, and $r=28$. According to [12], there exists an invariant ellipsoid $\mathcal{E}$ centered at $(0,0,2r)$ that contains all the $\omega$-limit sets of the system, which include three equilibria and the Lorenz attractor. Let $\mathcal{X}=\mathcal{E}$. Since $\mathcal{X}$ is invariant and compact, all trajectories are precompact in $\mathcal{X}$. Thus, by Corollary 3, there does not exist a one-to-one linear immersion for the Lorenz system on $\mathcal{X}$.
@@ 4 Implication for Learning Linear Immersions from Data
Assumption 1.
The system $\dot{x}=f(x)$ satisfies conditions (T2) and (T3) of Theorem 1 on a path-connected forward-invariant subset $\mathcal{X}$ of $\mathcal{M}$.
Generally, given a fixed sampling time $\tau$, a set of $N$ pairs $\{(x_{l},x_{l}^{+})\}_{l=1}^{N}$ where $x_{l}^{+}=\varphi(\tau,x_{l})$ for all $l$, and an integer $m>0$, the task of learning linear immersions $F:\mathcal{X}\rightarrow\mathbb{R}^{m}$ involves finding the following set
$\mathcal{F}^{*}(\tau,N)=\{F^{*}\in\mathcal{C}(\mathcal{X},\mathbb{R}^{m})\mid\exists A^{*}\in\mathbb{R}^{m\times m},$
@@
where $\mathcal{C}(\mathcal{X},\mathbb{R}^{m})$ is the space of continuous functions from $\mathcal{X}$ to $\mathbb{R}^{m}$. The state pairs $\{(x_{l},x_{l}^{+})\}_{l=1}^{N}$ can be extracted from a single trajectory or multiple trajectories of the system. Essentially, $\mathcal{F}^{*}(\tau,N)$ is the set of continuous functions that satisfy the condition in (3) for linear immersions at finitely many points $\{x_{l}\}_{l=1}^{N}$ and a fixed time step $\tau$.
@@
$(F^{*},A^{*})\in\underset{\begin{subarray}{c}F\in\mathcal{C}(\mathcal{X},\mathbb{R}^{m}),\\ A\in\mathbb{R}^{m\times m}\end{subarray}}{\arg\min}$ $~{}\sum_{l=1}^{N}\|F(x_{l}^{\prime})-e^{A\tau}F(x_{l})\|$
@@
The set $\mathcal{F}^{*}(\tau,N)$ in (25) corresponds to the solutions $F^{*}$ of this problem that give zero objective value, i.e., those interpolating the data perfectly.
Theorem 1 shows that under Assumption 1, any linear immersion $F$ satisfies that for all pairs $\Omega_{i}$, $\Omega_{j}$ in $\mathcal{W}^{+}$,
@@
A critical question here is if any learned linear immersion in $\mathcal{F}^{*}(\tau,N)$ would also share this property in (27). As a sanity check, note that any constant function $F^{*}(\cdot)=C$ for some $C\in\mathcal{Z}$ belongs to $\mathcal{F}^{*}(\tau,N)$ (with the corresponding $A^{*}=0$) and does map every $\omega$-limit set to the same set. However, this may not hold for every learned linear immersion in $\mathcal{F}^{*}(\tau,N)$ since the learned linear immersion only satisfies (3) at finitely many points (finitely many constraints), while a true linear immersion needs to satisfy (3) everywhere in $\mathcal{X}$ at all times (uncountably many constraints). However, in Theorem 4.3, we show that any function that does not collapse all $\omega$-limit sets would be excluded from $\mathcal{F}^{*}(\tau,N)$ for small enough $\tau$ and large enough $N$. The following proposition provides a crucial step for this result.
Proposition 4.2.
Let $\{x_{l}\}_{l=1}^{\infty}$ be a dense subset of $\mathcal{X}$. Let $F$ be a continuous mapping from $\mathcal{X}$ to $\mathbb{R}^{m}$. If for all $t>0$, there exists a sampling time $\tau\in(0,t]$ such that $F\in\mathcal{F}^{*}(\tau,N)$ for all $N>1$, then $F$ is a linear immersion of the system.
@@
Suppose that a continuous function $F$ satisfies the conditions in Theorem 4.2. Then, for all $j\in\mathbb{N}$, there exists a sampling time $\tau_{j}\in(0,2^{-j}]$ such that $F\in\mathcal{F}^{*}(\tau_{j},N)$ for all $N\geq 1$. Clearly, the positive sequence $\tau_{j}$ converges to zero as $j$ goes to infinity. Also, for each $j$, $F\in\lim_{N\rightarrow\infty}\mathcal{F}^{*}(\tau_{j},N)=:\mathcal{F}^{*}(\tau_{j})$. Note that the limit $\mathcal{F}^{*}(\tau_{j})$ is well defined since the set sequence $\mathcal{F}^{*}(\tau_{j},N)$ monotonically shrinks as $N$ increases.
@@
We pick an arbitrary $t>0$. Since $\tau_{j}$ is positive and converging to zero (recall that $\tau_{j}\in(0,2^{-j}]$), there exists $k_{j}\in\mathbb{N}$ such that the sequence $\bar{t}_{j}:=\sum_{i=1}^{j}k_{j}\cdot\tau_{j}$ converges to $t$ as $j$ goes to infinity. Note that for all $j$, (29) implies that
@@
Now, we state the main result of this section.
Theorem 4.3.
Let $\{x_{l}\}_{l=1}^{\infty}$ be a dense subset of $\mathcal{X}$. Let $F$ be a continuous function such that $F(\Omega_{1})\not=F(\Omega_{2})$ for some $\Omega_{1}$ and $\Omega_{2}\in\mathcal{W}^{+}$. Then, there exists $t^{*}>0$ such that for each sampling time $\tau\in(0,t^{*}]$, there exists an $N>0$ such that $F\not\in\mathcal{F}^{*}(\tau,N)$.
@@
Let $F$ be a continuous function that does not map every $\Omega\in\mathcal{W}^{+}$ to the same subset of $\mathbb{R}^{m}$. According to Theorem 1, since (T2) and (T3) hold, $F$ is not a linear immersion of the system. By the contrapositive of Proposition 4.2, for any $F$ not a linear immersion, there exists $t>0$ such that for all $\tau\in(0,t]$ and for some large enough $N$, $F$ is not in $\mathcal{F}^{*}(\tau,N)$. $\Box$
Theorem 4.3 reveals that any immersion candidate $F$ that can distinguish at least two $\omega$-limit sets in $\mathcal{X}$ would always be ruled out from $\mathcal{F}(\tau,N)$ by a small enough sampling time $\tau$ and a large enough sample size $N$. This is particularly the case for common algorithms that learn Koopman embeddings using a continuous parameterization, such as polynomials [35] and deep neural networks [37]. Hence, these algorithms will suffer from the issues identified in this section as long as they try to minimize the cost in (26) and achieve zero fitting error.
Remark 4.4.
@@
$\limsup_{j\rightarrow\infty}\lim_{N\rightarrow\infty}\mathcal{F}^{*}(\tau_{j},N)\subseteq\mathcal{C}(\mathcal{W}^{+}),$ (32)
where $\mathcal{C}(\mathcal{W}^{+})$ is the set of functions $F\in\mathcal{C}(\mathcal{X},\mathbb{R}^{m})$ such that $F(\Omega_{1})=F(\Omega_{2})$ for any $\Omega_{1}$ and $\Omega_{2}\in\mathcal{W}^{+}$.
Remark 4.5.